OpenStudy (anonymous):

Let X~(B6,0.35), find P(X<4|X>2)

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

anyone help please?? I really need to know it for my exams tomorrow D:

OpenStudy (anonymous):

What kind of problem is this

OpenStudy (anonymous):

stats @optiquest

OpenStudy (misty1212):

i think i might have an idea but not sure the question is what it the probability of less that 4 successes given that there are greater than 2 right?

OpenStudy (misty1212):

are you using a table of some kind?

OpenStudy (anonymous):

@misty1212 yes

OpenStudy (anonymous):

there is no table provided but I made my own

jimthompson5910 (jim_thompson5910):

are you allowed to use a calculator? if so, which type do you have?

OpenStudy (anonymous):

gdc ti-84 plus; yes you can use calc @jim_thompson5910

jimthompson5910 (jim_thompson5910):

I don't know what the gdc stands for, but I do have a TI84 plus you can use the TI84 plus to find binomial probabilities quickly hit the "2nd" key, then hit the "vars" key then scroll down to "binomcdf" the binomcdf function has this basic template binomcdf(trials, probability, value) http://tibasicdev.wikidot.com/binomcdf so you'll type in 6 for the trials, then 0.35 for the probability. The "value" is the value k such that \(\Large P(B \le k) = m\) is true. The value of m is the result of the binomcdf function

jimthompson5910 (jim_thompson5910):

Example: calculating P(X <= 4) binomcdf(6, 0.35, 4) = 0.977678

OpenStudy (anonymous):

@jim_thompson5910 I know how to do that, but for this particular question, P(X<4|X>2), what are the steps you have to take?

jimthompson5910 (jim_thompson5910):

first you need to find P(X > 2) use this idea \[\Large P(X \le 2) + P(X > 2) = 1\] \[\Large P(X > 2) = 1 - P(X \le 2)\]

jimthompson5910 (jim_thompson5910):

to calculate \(\Large P(X \le 2)\) you would type in binomcdf(6, 0.35, 2)

OpenStudy (anonymous):

Yep I did that. for this question, I tried doing P(X>2)/P(X<4) but its not correct...?

jimthompson5910 (jim_thompson5910):

Conditional Probability \[\Large P(A | B) = \frac{P(\text{A and B})}{P(B)}\]

jimthompson5910 (jim_thompson5910):

A = event that X < 4 B = event that X > 2

OpenStudy (anonymous):

so its P(X>2) x P(X<4) all divided by P(X<4)?

OpenStudy (anonymous):

won't it just be P(X>2)

jimthompson5910 (jim_thompson5910):

P(A and B) = P(X < 4 and X > 2) = P(X = 3) = ??

jimthompson5910 (jim_thompson5910):

you use the binompdf for single binomial probabilities (as opposed to binomcdf which adds up a bunch of probabilities)

OpenStudy (anonymous):

@jim_thompson5910 what do you mean?

jimthompson5910 (jim_thompson5910):

what is the probability that X = 3?

OpenStudy (anonymous):

0.8825.... but why do you need this number?

jimthompson5910 (jim_thompson5910):

no you don't type in binomcdf(6, 0.35, 3) for this part you type in binompdf(6, 0.35, 3)

OpenStudy (anonymous):

ohhh oops. so its 0.2354

OpenStudy (anonymous):

ahh i get it now!! the answer is 0.667. thank you so much!

jimthompson5910 (jim_thompson5910):

A = event that X < 4 B = event that X > 2 \[\Large P(A | B) = \frac{P(\text{A and B})}{P(B)}\] \[\Large P(X < 4 | X > 2) = \frac{P(X < 4 \text{ and } X > 2)}{P(X > 2)}\] \[\Large P(X < 4 | X > 2) = \frac{P(X = 3)}{P(X > 2)}\] \[\Large P(X < 4 | X > 2) = \frac{0.2354909375}{0.35291484}\] \[\Large P(X < 4 | X > 2) = \frac{0.2354909375}{0.35291484}\] \[\Large P(X < 4 | X > 2) = 0.66727411491112\]

jimthompson5910 (jim_thompson5910):

yes correct

OpenStudy (anonymous):

@jim_thompson5910 thank you so so so much!! would you be able to help me in another question?

jimthompson5910 (jim_thompson5910):

sure